# Time-Dependent Black Hole Lensing and the Unified Weak-to-Strong Deflection Framework

### v1

@@ III.2 Far-field (Born) expansion and centroid wobble
$+\mathcal{O}\,\left(\frac{M^{2}}{b^{2}},\varepsilon\frac{M}{b^{2}}\right),$ (30)
where $\hat{\bm{e}}_{b}$ is the outward radial unit vector in the $(X,Y)$ plane and $\bm{\mathcal{A}}=(c_{X}\mathcal{F}_{X},c_{Y}\mathcal{F}_{Y})$. That is, on the observer’s screen we use the orthonormal basis $\{\hat{\bm{e}}_{X},\hat{\bm{e}}_{Y}\}$ and define the screen radial unit vector $\hat{\bm{e}}_{b}\equiv(X\,\hat{\bm{e}}_{X}+Y\,\hat{\bm{e}}_{Y})/b$. The small $\mathcal{O}(M/b^{2})$ terms arise from curvature corrections to Eq. (26) and from subleading pieces of the kernel in Eq. (20).
Figure 2: Far-field, QNM-driven deflection amplitude vs impact parameter. The $\mathcal{O}(\varepsilon)$ correction from the Born map ($\omega_{I}<0$) behaves as $|\delta\bm{\alpha}(b,t_{o})|\sim|c_{A}(\omega)\,\mathcal{F}_{A}(\omega b)|/b$, leading to the $1/b$ falloff summarized in Eq. (30). Mild oscillations from the Fresnel factor $\mathcal{F}_{A}$ encode the finite-frequency sampling of the QNM along the nearly straight path, Eq. (26).
As seen in Fig. 2, the $\mathcal{O}(\varepsilon)$ deflection decays like $1/b$, with gentle structure governed by $\mathcal{F}_{A}(\omega b)$. This is the weak-field imprint that, when averaged over a bundle, yields the centroid wobble in Eq. (32).
To connect with observables we project Eq. (30) onto the screen angles $(\theta_{X},\theta_{Y})=(X/r_{o},Y/r_{o})$ defined in Eq. (13). For a narrow bundle of rays of fixed $b$ arriving at $(X,Y)$, the apparent position shift is
$\Delta\bm{\theta}(b,t_{o})=\frac{\bm{\alpha}(b,t_{o})}{r_{o}}$ (31)
up to the usual thin-deflection identification valid at $r_{o}\gg M$. We define the image centroid $\bm{\theta}_{c}(t_{o})$ as the intensity-weighted average position of the bundle on the screen; within the present purely theoretical treatment, we capture its dynamics by the average of Eq. (31) over the bundle’s narrow impact-parameter distribution. The centroid, therefore, exhibits a harmonic “wobble” at the QNM frequency:
@@
Figure 3: Centroid wobble in the weak field from Eq. (32). The components $\theta_{c,X}(t_{o})$ and $\theta_{c,Y}(t_{o})$ execute a damped harmonic motion at the QNM frequency $\omega=\omega_{R}+i\omega_{I}$ with $\omega_{I}<0$, with complex amplitude set by $\mathcal{B}(\omega)$. This panel visualizes the $\mathcal{O}(\varepsilon)$, gauge-safe centroid modulation predicted by the Born expansion in Eqs. (27)-(30).
As seen in Fig. 3, the centroid motion is a clean, single-frequency probe of the perturbation, with decay governed by $\omega_{I}<0$. The phase is inherited from the line-of-sight kernel in Eq. (24), anticipating the phase-locked relations with strong-field observables established in Sec. V. Eq. (32) provides the weak-field imaging diagnostic used later to match continuously to the strong-deflection modulations derived from the same master integral in Sec. IV.
@@ IV Time-Dependent Strong-Deflection Lensing
We now extract from the master Born representation in Eqs. (19)-(24) the strong-deflection limit $SDL$ relevant to trajectories that skim the photon sphere. The analysis is controlled by the small parameter $\Delta\equiv b/b_{c}-1$ with $0<\Delta\ll 1$, where $b$ is the impact parameter defined by Eq. (13) and $b_{c}$ denotes the critical value associated with the unstable circular null orbit. The time dependence enters exclusively through the $(\ell,m)=(2,0)$ QNM perturbation of Sec. II, which modulates the SDL coefficients and, in particular, the critical scale $b_{c}$. Our goal in this section is to obtain a time-dependent generalization of the logarithmic SDL law from the same kernel that produced the Born expansion in Sec. III.
@@ IV.1 Near-photon-sphere expansion
On the Schwarzschild background, the circular photon orbit sits at the photon-sphere radius Chandrasekhar (1985)
$r_{\mathrm{ph}}=3M,\qquad b_{c}=\frac{r_{\mathrm{ph}}}{\sqrt{f(r_{\mathrm{ph}})}}=3\sqrt{3}M.$ (33)
For $b\to b_{c}^{+}$ the bending angle $\alpha_{0}(b)$ obtained from Eq. (17) exhibits the well-known logarithmic divergence, Bozza (2002); Tsukamoto (2017)
$\alpha_{0}(b)=-\bar{a}\,\ln\,\left(\frac{b}{b_{c}}-1\right)+\bar{b}+\mathcal{O}\,\left((b-b_{c})\ln|b-b_{c}|\right),$ (34)
with the Schwarzschild values $\bar{a}=1$, and $\bar{b}=\ln\,\left[216(7-4\sqrt{3})\right]-\pi$ if one adopts the standard normalization. The constant $\bar{b}$ depends on the subtraction convention for the $\pi$ in the strong-deflection expansion; we follow the standard normalization commonly used for Schwarzschild. Eq. (34) follows by expanding the integral in Eq. (17) around the double root of the radial potential at $r_{\mathrm{ph}}$.
We now switch on the perturbation of Eq. (3) and use the time-dependent Born representation, Eqs. (19)-(24), to determine how the coefficients in Eq. (34) are modulated. Physically, the integral in Eq. (24) becomes dominated by the segment of the unperturbed trajectory that executes a long “whirl” near $r_{\mathrm{ph}}$, where small changes in the local geometry feed into large changes of the total deflection. It is therefore convenient to introduce the near-photon-sphere coordinate
$\rho\equiv\frac{r-r_{\mathrm{ph}}}{r_{\mathrm{ph}}},\qquad|\rho|\ll 1,$ (35)
and to parameterize the background null ray by the angle $\varphi$ accumulated during the whirl phase. In this regime the unperturbed relations implied by Eq. (17) reduce to the standard form
$\varphi(\rho)\simeq-\frac{1}{\sqrt{\kappa}}\,\ln|\rho|+\varphi_{0},\qquad\kappa=\frac{1}{2}r_{\mathrm{ph}}^{2}\,\left(\partial_{r}^{2}V_{\mathrm{eff}}\right)_{r_{\mathrm{ph}}},$ (36)
where $V_{\mathrm{eff}}$ is the Schwarzschild null effective potential ($V_{\rm eff}=f\,L^{2}/r^{2}$ for equatorial motion), and $\kappa>0$ encodes the instability exponent of the circular orbit. The logarithm in Eq. (36) is the origin of the SDL behavior in Eq. (34).
To first order in the perturbation amplitude $\varepsilon$, the metric perturbation $h_{\mu\nu}$ sampled along the whirl contributes to the kernel in Eq. (24) through $\partial_{\mu}h_{\alpha\beta}p_{0}^{\alpha}p_{0}^{\beta}$ evaluated at $r\simeq r_{\mathrm{ph}}$. Exploiting Eq. (36) to trade the affine parameter for $\varphi$, the correction $\delta\bm{\alpha}$ in Eq. (19) can be written, up to terms that remain finite as $b\to b_{c}^{+}$, as
@@
$+\mathcal{O}\,\left((b-b_{c})\ln|b-b_{c}|\right),$ (37)
@@
Combining Eqs. (34) and (37) with Eq. (19), we arrive at a compact time-dependent SDL law of the form
@@
$+\mathcal{O}\,\left((b-b_{c})\ln|b-b_{c}|\right),$ (38)
@@
$b_{c}(t_{o})$ $=b_{c}\left[1+\varepsilon\,\mathrm{Re}\left\{e^{-i\omega t_{o}}\,\beta_{1}(\omega)\right\}\right],$ (39)
with $a_{1},b_{1},\beta_{1}$ complex response coefficients that follow from the same near-photon-sphere kernel that produced $\mathcal{C}_{A},\mathcal{D}_{A}$ in Eq. (37). Eq. (38) is nothing but Eq. (34) with time-dependent coefficients. The modulation of $b_{c}$ in Eq. (39) can be interpreted geometrically as a small shift of the photon-sphere location,
$r_{\mathrm{ph}}(t_{o})=3M+\varepsilon\,\mathrm{Re}\left\{e^{-i\omega t_{o}}\,\delta r_{\mathrm{ph}}(\omega)\right\},$ (40)
propagated to $b_{c}(t_{o})$ through the algebraic relation in Eq. (33). The quantity $\delta r_{\mathrm{ph}}$ is determined by the $\mathcal{O}(\varepsilon)$ change of the null effective potential extremum in the perturbed geometry and is therefore a gauge-invariant statement at the order of interest, given our use of the Zerilli master field.
It is often convenient to linearize Eq. (38) explicitly in $\varepsilon$. Writing $\Delta=b/b_{c}-1$ and keeping only $\mathcal{O}(\varepsilon)$ terms, we obtain
@@
$+\bar{a}\frac{\beta_{1}(\omega)}{\Delta}+b_{1}(\omega)\biggr]\biggr\}+\mathcal{O}\,\left(\varepsilon\,\delta\ln\Delta\right),$ (41)
where $\alpha_{0}(b)$ is given by Eq. (34), and $\bar{a}=1$ for the Schwarzschild case. By axisymmetry, the leading $O(\varepsilon)$ near-critical deflection is radial on the screen; we therefore drop tangential components in what follows. The term proportional to $\beta_{1}/\Delta$ originates from expanding $\ln\,\left(b/b_{c}(t_{o})-1\right)$ and makes explicit that an oscillatory shift of $b_{c}$ produces the largest time-dependent contribution as $b\to b_{c}^{+}$. With our $P_{2}$ normalization, $\beta_{1}(\omega)=\frac{1}{4}\,H_{0}(r_{\mathrm{ph}};\omega)$.
Figure 4: Near-critical deflection from the time-dependent SDL law. Shown is $\alpha(b,t_{o})$ versus $\ln\Delta$ with $\Delta=b/b_{c}-1\ll 1$, as predicted by Eq. (38) and its linearized form Eq. (41). The dominant $1/\Delta$ sensitivity arises from the modulation of $b_{c}(t_{o})$ through $\beta_{1}(\omega)$, while the slope receives a small time-dependent correction from $a_{1}(\omega)$. Parallel, slightly offset curves at different $t_{o}$ illustrate the coherent QNM-driven modulation.
Figure 4 makes explicit the near-critical amplification: the $\beta_{1}/\Delta$ piece dominates as $\Delta\to 0^{+}$, while the $a_{1}\ln\Delta$ term induces a mild slope variation. This behavior feeds directly into the image hierarchy and photon-ring modulations derived in Sec. IV.2, and it will be matched to the weak-field kernel in Sec. V.1. Eq. (41) is the central near-critical result that will feed, in Sec. IV.2, the logarithmic time delay, the multiplicity of relativistic images, and their photon-ring modulations, all sharing the common QNM frequency inherited from Eq. (24).
@@ IV.2 Time delays, multiplicity, and photon-ring modulation
Eq. (38) implies that, for each winding number $n\in\mathbb{N}$, relativistic images arise from solutions of the lens equation with total bending $\alpha(b,t_{o})\approx 2\pi n+\Delta\alpha_{n}$ where $|\Delta\alpha_{n}|\ll 1$. Using Eq. (38) at fixed observer time $t_{o}$ and solving for $b$ near $b_{c}(t_{o})$ gives the instantaneous sequence of critical impact parameters
@@
$s_{n}(t_{o})$ $\equiv\exp\,\left(\frac{\bar{b}(t_{o})-2\pi n}{\bar{a}(t_{o})}\right),$ (42)
@@
To expose the imprint of the $(\ell,m)=(2,0)$ QNM, we linearize Eq. (42) in $\varepsilon$ using Eq. (39). Writing $\theta_{c}\equiv b_{c}/r_{o}$ and $s_{n}^{(0)}\equiv\exp\,\left[(\bar{b}-2\pi n)/\bar{a}\right]$, we obtain
@@
$\delta\theta_{n}=\theta_{c}\,\beta_{1}\left(1+s_{n}^{(0)}\right)+\theta_{c}\,s_{n}^{(0)}\,\left(\frac{b_{1}}{\bar{a}}-\frac{\bar{b}-2\pi n}{\bar{a}^{2}}\,a_{1}\right).$ (44)
Eqs. (43)-(44) show that all relativistic images wobble coherently at the QNM frequency $\omega$, with an amplitude that decays geometrically with $n$ through $s_{n}^{(0)}$.
Figure 5: Relativistic image hierarchy and photon ring. The angular positions $\theta_{n}(t_{o})$ are shown versus winding number $n$ for several snapshots in $t_{o}$, with the instantaneous ring radius $\theta_{\rm ring}(t_{o})=b_{c}(t_{o})/r_{o}$ indicated. The geometric approach $\theta_{n}\to\theta_{\rm ring}$ is governed by $s_{n}^{(0)}=\exp[(\bar{b}-2\pi n)/\bar{a}]$, while the coherent, small-amplitude wobble of all $\theta_{n}$ is driven by the QNM via the response amplitudes in Eq. (44).
Figure 5 displays the expected geometric accumulation of images toward the ring, with a phase-locked modulation at frequency $\omega$ inherited from Eq. (24). The $\beta_{1}$ contribution shifts the ring itself, while the $a_{1}$ and $b_{1}$ terms imprint the subleading $n$-dependent offsets in Eq. (44).
The same near-photon-sphere control that yields Eq. (38) provides the coordinate travel time for a near-critical ray. Denote by $T(b,t_{o})$ the total coordinate time elapsed between emission and detection for a ray of impact parameter $b$. By standard SDL analysis, one finds a logarithmic form
@@
$+\tilde{b}(t_{o})+n\,T_{\mathrm{whirl}}(t_{o})+\ldots,$ (45)
where $n$ counts the net azimuthal winds accumulated during the whirl phase, $T_{\mathrm{whirl}}(t_{o})$ is the period of one circular photon orbit measured in Schwarzschild time, and $\tilde{a}(t_{o}),\tilde{b}(t_{o})$ are time-dependent analogues of $\bar{a}(t_{o}),\bar{b}(t_{o})$ governed by the same Born kernel. For Schwarzschild, in the static limit, we have Chandrasekhar (1985)
$T_{\mathrm{whirl}}=\frac{2\pi}{\Omega_{\mathrm{ph}}},\qquad\Omega_{\mathrm{ph}}=\frac{1}{3\sqrt{3}M},$ (46)
so that successive relativistic images are separated, to leading order, by $2\pi/\Omega_{\mathrm{ph}}$ in arrival time. Allowing for $\mathcal{O}(\varepsilon)$ modulations analogous to Eq. (39), we write
@@
$T_{\mathrm{whirl}}(t_{o})$ $=T_{\mathrm{whirl}}\left[1+\varepsilon\,\mathrm{Re}\left\{e^{-i\omega t_{o}}\,\gamma_{1}\right\}\right],$ (47)
with complex response coefficients $\tilde{a}_{1},\tilde{b}_{1},\gamma_{1}$ fixed by the same near-photon-sphere kernel that determines $a_{1},b_{1},\beta_{1}$ in Eq. (39). Evaluating Eq. (45) at $b=b_{n}(t_{o})$ and forming differences, the inter-image delay becomes

### v2

@@ III.2 Far-field (Born) expansion and centroid wobble
$+\mathcal{O}\,\left(\frac{M^{2}}{b^{2}},\varepsilon\frac{M}{b^{2}}\right),$ (III.13)
where $\hat{\bm{e}}_{b}$ is the outward radial unit vector in the $(X,Y)$ plane and $\bm{\mathcal{A}}=(c_{X}\mathcal{F}_{X},c_{Y}\mathcal{F}_{Y})$. That is, on the observer’s screen we use the orthonormal basis $\{\hat{\bm{e}}_{X},\hat{\bm{e}}_{Y}\}$ and define the screen radial unit vector $\hat{\bm{e}}_{b}\equiv(X\,\hat{\bm{e}}_{X}+Y\,\hat{\bm{e}}_{Y})/b$. The small $\mathcal{O}(M/b^{2})$ terms arise from curvature corrections to Eq. (III.9) and from subleading pieces of the kernel in Eq. (III.4).
Figure 2: Far-field, QNM-driven deflection amplitude vs impact parameter. The $\mathcal{O}(\varepsilon)$ Born correction obeys $|\delta\alpha(b,t_{o})|\sim|A(b,\omega)|/b$ with mild oscillations from the Fresnel factor $F_{A}(\omega b)$ [Eqs. (III.10)-(III.13)]. Using logarithmic axes makes the $1/b$ decay explicit (slope 1-1). A dashed curve shows a reference $\propto 1/b$ law scaled at a pivot $b_{\rm ref}$ for comparison.
As seen in Fig. 2, the $\mathcal{O}(\varepsilon)$ deflection decays like $1/b$, with gentle structure governed by $\mathcal{F}_{A}(\omega b)$. This is the weak-field imprint that, when averaged over a bundle, yields the centroid wobble in Eq. (III.15).
To connect with observables we project Eq. (III.13) onto the screen angles $(\theta_{X},\theta_{Y})=(X/r_{o},Y/r_{o})$ defined in Eq. (II.13). For a narrow bundle of rays of fixed $b$ arriving at $(X,Y)$, the apparent position shift is
$\Delta\bm{\theta}(b,t_{o})=\frac{\bm{\alpha}(b,t_{o})}{r_{o}}$ (III.14)
up to the usual thin-deflection identification valid at $r_{o}\gg M$. We define the image centroid $\bm{\theta}_{c}(t_{o})$ as the intensity-weighted average position of the bundle on the screen; within the present purely theoretical treatment, we capture its dynamics by the average of Eq. (III.14) over the bundle’s narrow impact-parameter distribution. The centroid, therefore, exhibits a harmonic wobble at the QNM frequency:
@@
Figure 3: Centroid wobble in the weak field from Eq. (III.15). The components $\theta_{c,X}(t_{o})$ and $\theta_{c,Y}(t_{o})$ execute a damped harmonic motion at the QNM frequency $\omega=\omega_{R}+i\omega_{I}$ with $\omega_{I}<0$, with complex amplitude set by $\mathcal{B}(\omega)$. This panel visualizes the $\mathcal{O}(\varepsilon)$, gauge-safe centroid modulation predicted by the Born expansion in Eqs. (III.10)-(III.13).
As seen in Fig. 3, the centroid motion is a clean, single-frequency probe of the perturbation, with decay governed by $\omega_{I}<0$. The phase is inherited from the line-of-sight kernel in Eq. (III.7), anticipating the phase-locked relations with strong-field observables established in Sec. V. Eq. (III.15) provides the weak-field imaging diagnostic used later to match continuously to the strong-deflection modulations derived from the same master integral in Sec. IV.
@@ IV Time-Dependent Strong-Deflection Lensing
We now extract from the master Born representation in Eqs. (III.3)-(III.7) the strong-deflection limit $SDL$ relevant to trajectories that skim the photon sphere. The analysis is controlled by the small parameter $\Delta\equiv b/b_{c}-1$ with $0<\Delta\ll 1$, where $b$ is the impact parameter defined by Eq. (II.13) and $b_{c}$ denotes the critical value associated with the unstable circular null orbit. The time dependence enters exclusively through the $(\ell,m)=(2,0)$ QNM perturbation of Sec. II, which modulates the SDL coefficients and, in particular, the critical scale $b_{c}$. Our goal in this section is to obtain a time-dependent generalization of the logarithmic SDL law from the same kernel that produced the Born expansion in Sec. III.
@@ IV.1 Near-photon-sphere expansion
On the Schwarzschild background, the circular photon orbit sits at the photon-sphere radius [13, 14]
$r_{\mathrm{ph}}=3M,\qquad b_{c}=\frac{r_{\mathrm{ph}}}{\sqrt{f(r_{\mathrm{ph}})}}=3\sqrt{3}M.$ (IV.1)
For $b\to b_{c}^{+}$ the bending angle $\alpha_{0}(b)$ obtained from Eq. (III.1) exhibits the well-known logarithmic divergence, [9, 41]
$\alpha_{0}(b)=-\bar{a}\,\ln\,\left(\frac{b}{b_{c}}-1\right)+\bar{b}+\mathcal{O}\,\left((b-b_{c})\ln|b-b_{c}|\right),$ (IV.2)
with the Schwarzschild values $\bar{a}=1$, and $\bar{b}=\ln\,\left[216(7-4\sqrt{3})\right]-\pi$ if one adopts the standard normalization. The constant $\bar{b}$ depends on the subtraction convention for the $\pi$ in the strong-deflection expansion; we follow the standard normalization commonly used for Schwarzschild. Eq. (IV.2) follows by expanding the integral in Eq. (III.1) around the double root of the radial potential at $r_{\mathrm{ph}}$.
We now switch on the perturbation of Eq. (II.3) and use the time-dependent Born representation, Eqs. (III.3)-(III.7), to determine how the coefficients in Eq. (IV.2) are modulated. Physically, the integral in Eq. (III.7) becomes dominated by the segment of the unperturbed trajectory that executes a long whirl near $r_{\mathrm{ph}}$, where small changes in the local geometry feed into large changes of the total deflection. It is therefore convenient to introduce the near-photon-sphere coordinate
$\rho\equiv\frac{r-r_{\mathrm{ph}}}{r_{\mathrm{ph}}},\qquad|\rho|\ll 1,$ (IV.3)
and to parameterize the background null ray by the angle $\varphi$ accumulated during the whirl phase. In this regime the unperturbed relations implied by Eq. (III.1) reduce to the standard form
$\varphi(\rho)\simeq-\frac{1}{\sqrt{\kappa}}\,\ln|\rho|+\varphi_{0},\qquad\kappa=\frac{1}{2}r_{\mathrm{ph}}^{2}\,\left(\partial_{r}^{2}V_{\mathrm{eff}}\right)_{r_{\mathrm{ph}}},$ (IV.4)
where $V_{\mathrm{eff}}$ is the Schwarzschild null effective potential ($V_{\rm eff}=f\,L^{2}/r^{2}$ for equatorial motion), and $\kappa>0$ encodes the instability exponent of the circular orbit. The logarithm in Eq. (IV.4) is the origin of the SDL behavior in Eq. (IV.2).
To first order in the perturbation amplitude $\varepsilon$, the metric perturbation $h_{\mu\nu}$ sampled along the whirl contributes to the kernel in Eq. (III.7) through $\partial_{\mu}h_{\alpha\beta}p_{0}^{\alpha}p_{0}^{\beta}$ evaluated at $r\simeq r_{\mathrm{ph}}$. Exploiting Eq. (IV.4) to trade the affine parameter for $\varphi$, the correction $\delta\bm{\alpha}$ in Eq. (III.3) can be written, up to terms that remain finite as $b\to b_{c}^{+}$, as
@@
$+\mathcal{O}\,\left((b-b_{c})\ln|b-b_{c}|\right),$ (IV.5)
@@
Combining Eqs. (IV.2) and (IV.5) with Eq. (III.3), we arrive at a compact time-dependent SDL law of the form
@@
$+\mathcal{O}\,\left((b-b_{c})\ln|b-b_{c}|\right),$ (IV.6)
@@
$b_{c}(t_{o})$ $=b_{c}\left[1+\varepsilon\,\mathrm{Re}\left\{e^{-i\omega t_{o}}\,\beta_{1}(\omega)\right\}\right],$ (IV.7)
with $a_{1},b_{1},\beta_{1}$ complex response coefficients that follow from the same near-photon-sphere kernel that produced $\mathcal{C}_{A},\mathcal{D}_{A}$ in Eq. (IV.5). Eq. (IV.6) is nothing but Eq. (IV.2) with time-dependent coefficients. The modulation of $b_{c}$ in Eq. (IV.7) can be interpreted geometrically as a small shift of the photon-sphere location,
$r_{\mathrm{ph}}(t_{o})=3M+\varepsilon\,\mathrm{Re}\left\{e^{-i\omega t_{o}}\,\delta r_{\mathrm{ph}}(\omega)\right\},$ (IV.8)
propagated to $b_{c}(t_{o})$ through the algebraic relation in Eq. (IV.1). The quantity $\delta r_{\mathrm{ph}}$ is determined by the $\mathcal{O}(\varepsilon)$ change of the null effective potential extremum in the perturbed geometry and is therefore a gauge-invariant statement at the order of interest, given our use of the Zerilli master field.
It is often convenient to linearize Eq. (IV.6) explicitly in $\varepsilon$. Writing $\Delta=b/b_{c}-1$ and keeping only $\mathcal{O}(\varepsilon)$ terms, we obtain
@@
$+\bar{a}\frac{\beta_{1}(\omega)}{\Delta}+b_{1}(\omega)\biggr]\biggr\}+\mathcal{O}\,\left(\varepsilon\,\Delta\ln\Delta\right)+\mathcal{O}\,\left(\varepsilon^{2}\right),$ (IV.9)
where $\alpha_{0}(b)$ is given by Eq. (IV.2), and $\bar{a}=1$ for the Schwarzschild case. By axisymmetry, the leading $O(\varepsilon)$ near-critical deflection is radial on the screen; we therefore drop tangential components in what follows. The term proportional to $\beta_{1}/\Delta$ originates from expanding $\ln\,\left(b/b_{c}(t_{o})-1\right)$ and makes explicit that an oscillatory shift of $b_{c}$ produces the largest time-dependent contribution as $b\to b_{c}^{+}$. With our $P_{2}$ normalization, $\beta_{1}(\omega)=\frac{1}{4}\,H_{0}(r_{\mathrm{ph}};\omega)$.
Our first-order, line-of-sight (Born) expansion keeps the Schwarzschild background deflection $\alpha_{0}(b)$ exact and expands in the QNM amplitude $\varepsilon$. In the near-critical regime $\Delta\equiv b/b_{c}-1\ll 1$, Eq. (IV.9) shows that the dominant time-dependent correction scales as $\delta\alpha\sim\varepsilon\,\beta_{1}/\Delta$, whereas the background scales as $\alpha_{0}\sim-\ln\Delta$ [cf. Eq. (IV.2)]. A conservative perturbative criterion is therefore
$\mathcal{R}(b)\;\equiv\;\frac{|\varepsilon\,\beta_{1}|}{\Delta\,|\ln\Delta|}\ll 1\quad\Rightarrow\quad\Delta\gg\frac{|\varepsilon\,\beta_{1}|}{|\ln\Delta|}\,$, (IV.10)
which, up to the slowly varying logarithm, is the order-of-magnitude condition $\Delta\gtrsim\mathcal{O}(\varepsilon|\beta_{1}|)$. Equivalently, using the static SDL spacing $s_{n}^{(0)}=\exp[(\bar{b}-2\pi n)/\bar{a}]$ [Eq. (IV.12) with $\bar{a}=1$], the $n$-th relativistic image remains within the linear regime provided
$s_{n}^{(0)}\gg|\varepsilon\,\beta_{1}|\quad\Longleftrightarrow\quad n\lesssim\frac{1}{2\pi}\!\left[\ln\!\frac{1}{\varepsilon|\beta_{1}|}+\mathcal{O}(1)\right].$ (IV.11)
In words: the instantaneous ring-radius wobble $\delta\theta_{\rm ring}/\theta_{c}=\varepsilon\,|\beta_{1}|$ [Eq. (IV.19)] must be smaller than the image’s offset from the ring $\theta_{n}-\theta_{\rm ring}\simeq\theta_{c}\,s_{n}^{(0)}$.
Figure 4: Near-critical deflection from the time-dependent SDL law. Shown is $\alpha(b,t_{o})$ versus $\ln\Delta$ with $\Delta=b/b_{c}-1\ll 1$, as predicted by Eq. (IV.6) and its linearized form Eq. (IV.9). The dominant $1/\Delta$ sensitivity arises from the modulation of $b_{c}(t_{o})$ through $\beta_{1}(\omega)$, while the slope receives a small time-dependent correction from $a_{1}(\omega)$. Parallel, slightly offset curves at different $t_{o}$ illustrate the coherent QNM-driven modulation.
Figure 4 makes explicit the near-critical amplification: the $\beta_{1}/\Delta$ piece dominates as $\Delta\to 0^{+}$, while the $a_{1}\ln\Delta$ term induces a mild slope variation. This behavior feeds directly into the image hierarchy and photon-ring modulations derived in Sec. IV.2, and it will be matched to the weak-field kernel in Sec. V.1. Eq. (IV.9) is the central near-critical result that will feed, in Sec. IV.2, the logarithmic time delay, the multiplicity of relativistic images, and their photon-ring modulations, all sharing the common QNM frequency inherited from Eq. (III.7).
@@ IV.2 Time delays, multiplicity, and photon-ring modulation
Eq. (IV.6) implies that, for each winding number $n\in\mathbb{N}$, relativistic images arise from solutions of the lens equation with total bending [9, 46, 51] $\alpha(b,t_{o})\approx 2\pi n+\Delta\alpha_{n}$ where $|\Delta\alpha_{n}|\ll 1$. Throughout this section, $n\in\mathbb{N}$ denotes the winding number, i.e. the image order: a solution with total bending $\alpha(b,t_{o})\simeq 2\pi n+\Delta\alpha_{n}$ corresponds to the $n$-th relativistic image formed by a photon that executes $n$ azimuthal winds near the photon sphere ($n=1$ for one whirl, $n=2$ for two whirls, etc.). Using Eq. (IV.6) at fixed observer time $t_{o}$ and solving for $b$ near $b_{c}(t_{o})$ gives the instantaneous sequence of critical impact parameters
@@
$s_{n}(t_{o})$ $\equiv\exp\,\left(\frac{\bar{b}(t_{o})-2\pi n}{\bar{a}(t_{o})}\right),$ (IV.12)
@@
To expose the imprint of the $(\ell,m)=(2,0)$ QNM, we linearize Eq. (IV.12) in $\varepsilon$ using Eq. (IV.7). Writing $\theta_{c}\equiv b_{c}/r_{o}$ and $s_{n}^{(0)}\equiv\exp\,\left[(\bar{b}-2\pi n)/\bar{a}\right]$, we obtain
@@
$\delta\theta_{n}=\theta_{c}\,\beta_{1}\left(1+s_{n}^{(0)}\right)+\theta_{c}\,s_{n}^{(0)}\,\left(\frac{b_{1}}{\bar{a}}-\frac{\bar{b}-2\pi n}{\bar{a}^{2}}\,a_{1}\right).$ (IV.14)
Eqs. (IV.13)-(IV.14) show that all relativistic images wobble coherently at the QNM frequency $\omega$, with an amplitude that decays geometrically with $n$ through $s_{n}^{(0)}$.
Figure 5: Markers: solid points show the angular positions of the $n$-th relativistic image, $\theta_{n}(t_{o})$; the black dashed line shows the instantaneous photon-ring radius $\theta_{\rm ring}(t_{o})=b_{c}(t_{o})/r_{o}$ [Eq. (IV.19)]. The rapid convergence $\theta_{n}\to\theta_{\rm ring}$ with increasing $n$ follows the static SDL spacing $s_{n}^{(0)}=\exp[(\bar{b}-2\pi n)/\bar{a}]$, while the coherent, small-amplitude wobble of all $\theta_{n}$ with $t_{o}$ is the QNM-driven modulation predicted by Eqs. (IV.13)-(IV.14). Inset (semilog $y$-axis): the difference $|\Delta\theta_{n}|\equiv|\theta_{n}-\theta_{\rm ring}|$ versus $n$ at a fixed $t_{o}$, highlighting the geometric decay $|\Delta\theta_{n}|\simeq\theta_{c}\,s_{n}^{(0)}$ with baseline slope $\ln q_{0}=-2\pi$ (Schwarzschild, $\bar{a}=1$) and a weak, frequency-locked modulation from Eq. (V.10).
Figure 5 displays the expected geometric accumulation of images toward the ring, with a phase-locked modulation at frequency $\omega$ inherited from Eq. (III.7). The $\beta_{1}$ contribution shifts the ring itself, while the $a_{1}$ and $b_{1}$ terms imprint the subleading $n$-dependent offsets in Eq. (IV.14).
The same near-photon-sphere control that yields Eq. (IV.6) provides the coordinate travel time for a near-critical ray. Denote by $T(b,t_{o})$ the total coordinate time elapsed between emission and detection for a ray of impact parameter $b$. By standard SDL analysis, one finds a logarithmic form
@@
$+\tilde{b}(t_{o})+n\,T_{\mathrm{whirl}}(t_{o})+\ldots,$ (IV.15)
where $n$ counts the net azimuthal winds accumulated during the whirl phase, $T_{\mathrm{whirl}}(t_{o})$ is the period of one circular photon orbit measured in Schwarzschild time, and $\tilde{a}(t_{o}),\tilde{b}(t_{o})$ are time-dependent analogues of $\bar{a}(t_{o}),\bar{b}(t_{o})$ governed by the same Born kernel. For Schwarzschild, in the static limit, we have [19, 10, 8]
$T_{\mathrm{whirl}}=\frac{2\pi}{\Omega_{\mathrm{ph}}},\qquad\Omega_{\mathrm{ph}}=\frac{1}{3\sqrt{3}M},$ (IV.16)
so that successive relativistic images are separated, to leading order, by $2\pi/\Omega_{\mathrm{ph}}$ in arrival time. Allowing for $\mathcal{O}(\varepsilon)$ modulations analogous to Eq. (IV.7), we write
@@
$T_{\mathrm{whirl}}(t_{o})$ $=T_{\mathrm{whirl}}\left[1+\varepsilon\,\mathrm{Re}\left\{e^{-i\omega t_{o}}\,\gamma_{1}\right\}\right],$ (IV.17)
with complex response coefficients $\tilde{a}_{1},\tilde{b}_{1},\gamma_{1}$ fixed by the same near-photon-sphere kernel that determines $a_{1},b_{1},\beta_{1}$ in Eq. (IV.7). Evaluating Eq. (IV.15) at $b=b_{n}(t_{o})$ and forming differences, the inter-image delay becomes